OpenStudy (anonymous):

HHEELLPP !! !! !! What is the function for the arithmetic sequence 2, 7, 12, 17, 22, . . . . . ? A) A(n)=5+(n-1)(2) B) A(n)=2+(n-1)(6) C) A(n)=5x6 D) A(n)=2+(n-1)(5)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

u sure ?

OpenStudy (anonymous):

Plug in A(1) into each choice,

OpenStudy (anonymous):

a=2 c.d. d=7-2=5 \[A(n)=a+(n-1)d\]

OpenStudy (anonymous):

if a(1) = 2 and a(2) = 7 ... and so on then thats your answer

OpenStudy (anonymous):

so D ?